# A Longitudinal Study of Child Wellbeing Assessment via Online Interactions with a Social Robot

### v2

@@ 6 Discussion
Following the results of this study, and considering the future vision of integrating social robots into supporting mental health assessment, we propose a customised approach for a robot-led mental wellbeing assessments rather than a one-size-fits-all model. As such, the frequency of the evaluations should be decided in line with current guidelines  [^11] https://www.nhs.uk/mental-health/social-care-and-your-rights/mental-health-assessments/ that determine the severity of the conditions and the care provided. As seen in Figure 1 (b), robots’ multimodal sensing capabilities can be utilised to sense and identify potential wellbeing concerns in children through verbal and non-verbal cues. Child mental health practitioners can then review the observations collected by robots and make informed decisions. This approach provides genuine data and insights that enhance and support clinical decision-making, rather than merely guiding it. In addition, robots can also be programmed with "memory" abilities to learn from prior assessments for a particular child and focus on areas that might warrant further investigation. Robotised assessments can harness the time efficiency of ecological momentary assessment and intervention programs like Artemis-A  [^12] https://www.artemis-a.org/ which utilise computerised adaptive testing to significantly reduce the duration of follow-up assessments. This can alleviate the workload and pressure on existing psychological resources that rely solely on the verbal and written responses provided by children [32]. As such, robotised assessments can be used as valuable tools to (I) help children become more candid with their responses, (II) flag cases that should receive further attention from child mental health practitioners, (III) provide genuine data of verbal and non-verbal cues from authentic interactions that can support clinical decision making, and (IV) carry out time-efficient follow up assessments if needed, to develop a more comprehensive understanding of the children’s mental wellbeing over time and in ecological settings (e.g., public spaces, local General Practitioner’s surgery and schools). We aspire for robot-led assessments to serve as useful tools supporting the work of child mental health practitioners in diagnosing and creating treatment plans that monitor changes in the mental wellbeing of future generations in an ethical and responsible way.
7 conclusions and future works
Our work has unveiled compelling insights into longitudinally assessing children’s mental wellbeing through online interactions with a robot. We hope that robots can form a valuable addition to the mental wellbeing assessment procedure of children, to aid clinicians and psychologists in identifying cases with potential mental wellbeing concerns so that necessary clinical intervention can be provided in a timely manner. Although we have addressed the research questions defined at the outset of our work, our study has several limitations that we hope to address in the future. Firstly, we have only considered verbal responses from the participants. Future work will also focus on the non-verbal responses (e.g., facial expressions, audio) of the participants. Secondly, our research is a feasibility study conducted on a non-clinical population to determine whether and how SARs can be used as novel tools to contribute to the holistic understanding of mental health status. In future, we do hope to conduct case-controlled designs in clinical settings with clinical populations through clinician-administered measures. Finally, we have not taken into account wellbeing changes between online form-filling and the experiment sessions. In future, we aim to conduct self-report and parent-report questionnaires concurrently with the robot interaction session to avoid any potential fluctuations in mental wellbeing reporting.
@@ Acknowledgments
N. I. Abbasi is supported by the W.D. Armstrong Trust PhD Studentship and the Cambridge Trusts. G. Laban and H. Gunes are supported by the EPSRC project ARoEQ under grant ref. EP/R030782/1. T. Ford and P. Jones acknowledge support from the NIHR ARC East of England and the Cambridge BRC. The equipment used in the study was supported by the University of Cambridge’s Human Machine Collaboration (OHMC) Small Equipment Funding. The views expressed are those of the authors and not necessarily those of the NIHR or the Department of Health and Social Care.
@@
For the purpose of open access, the authors have applied a Creative Commons Attribution (CC BY) licence to any Author Accepted Manuscript version arising.

### v3

@@ 5. Discussion
Following the results of this study, and considering the future vision of integrating social robots into supporting mental health assessment, we propose a customised approach for a robot-led mental wellbeing assessments rather than a one-size-fits-all model. As such, the frequency of the evaluations should be decided in line with current guidelines  [^11] https://www.nhs.uk/mental-health/social-care-and-your-rights/mental-health-assessments/ that determine the severity of the conditions and the care provided. As seen in Figure 1 (b), robots’ multimodal sensing capabilities can be utilised to sense and identify potential wellbeing concerns in children through verbal and non-verbal cues. Child mental health practitioners can then review the observations collected by robots and make informed decisions. This approach provides genuine data and insights that enhance and support clinical decision-making, rather than merely guiding it. In addition, robots can also be programmed with ”memory” abilities to learn from prior assessments for a particular child and focus on areas that might warrant further investigation. Robotised assessments can harness the time efficiency of ecological momentary assessment and intervention programs like Artemis-A  [^12] https://www.artemis-a.org/ which utilise computerised adaptive testing to significantly reduce the duration of follow-up assessments. This can alleviate the workload and pressure on existing psychological resources that rely solely on the verbal and written responses provided by children (Laban et al., 2022a). As such, robotised assessments can be used as valuable tools to (I) help children become more candid with their responses, (II) flag cases that should receive further attention from child mental health practitioners, (III) provide genuine data of verbal and non-verbal cues from authentic interactions that can support clinical decision making, and (IV) carry out time-efficient follow up assessments if needed, to develop a more comprehensive understanding of the children’s mental wellbeing over time and in ecological settings (e.g., public spaces, local General Practitioner’s surgery and schools). We aspire for robot-led assessments to serve as useful tools supporting the work of child mental health practitioners in diagnosing and creating treatment plans that monitor changes in the mental wellbeing of future generations in an ethical and responsible way.
Finally, the use of adaptive robot behaviour in the context of mental well-being and HRI presents significant potential for enhancing personalised interactions, particularly through the integration of large language models (LLMs) (Zhang et al., 2023; Kim et al., 2024) and affective capabilities such as affect recognition from facial and vocal expressions, as well as behavioural cues (Spitale and Gunes, 2022; Spitale et al., 2024). By dynamically adjusting the robot’s responses based on users’ affective and psychological states, these adaptive systems can foster greater engagement, rapport, and deeper self-disclosure (Laban and Cross, 2024). However, to ensure accurate mental health assessments, these interactions should adhere to standard practices by following a structured, validated, and well-established assessment frameworks (Abbasi et al., 2024a). Hence, certain aspects of interaction need to remain structured and consistent in order to adhere to established protocols of wellbeing assessments. For instance, the RCADS and the SMFQ need to be conducted in accordance with standardised guidelines to maintain the validity and reliability of the findings (Angold et al., 1995; Sharp et al., 2006; Chorpita et al., 2015). The questionnaire statements could not be modified as doing so would compromise the validity of the psychological test under consideration (Stewart et al., 2012). As such, this highlights the need to balance flexibility with adherence to established protocols so as to ensure desired, valid and reliable outcomes (Abbasi et al., 2024a).
Consequently, the current adaptive approaches may not be particularly suitable for mental health assessments (see Axelsson et al., 2024, ”(R4) Verbal adaptation should be limited to preserve well-being practice efficacy”), especially with children. Moreover, despite the potential benefits of affective technologies to quality of interaction, there are critical ethical concerns related to privacy, data security, and the handling of sensitive emotional information. These adaptive capabilities rely on processing and sharing large amounts of personal data with third parties, raising concerns about the transparency of the algorithms and the potential for biased or inappropriate responses, especially in vulnerable populations like children (Kurian, 2024; Laban and Cross, 2024). Additionally, there are concerns regarding the potential for LLMs to inadvertently reinforce biases or generate inappropriate responses, which could be particularly harmful in sensitive mental health contexts (Spitale et al., 2024). For example, if the model produces responses misaligned with a child’s developmental needs or misinterprets the emotional nuances of a situation, it could negatively affect the child’s well-being (Kurian, 2023). The opaque nature of LLMs further complicates efforts to fully ensure the safety and appropriateness of these interactions, emphasising the need for real-time oversight and intervention (Liao and Vaughan, 2024; Harrer, 2023). Therefore, while adaptive LLM-driven behaviour holds promise for improving mental well-being outcomes in HRI, it is imperative to rigorously test for biases, ensure transparent data handling, and implement safeguards to prevent unintended consequences (e.g., see Tam et al., 2024). Future work should focus on refining these adaptive and affective capabilities to prioritise both the psychological benefits for users and the ethical management of their data, ensuring trust and safety in long-term HRI applications.
This calls for a further essential discussion addressing the ethical concerns surrounding the use of social robots in mental wellbeing assessments, particularly regarding trust, confidentiality, and managing children’s expectations of privacy. While social robots can foster trust and encourage self-disclosure due to their non-judgmental presence, it is crucial to communicate the limits of confidentiality clearly. Children and their guardians must understand that disclosures related to safety concerns may need to be shared with trusted adults, in line with safeguarding protocols. To balance trust with safeguarding duties, a tiered privacy management strategy could be implemented, ensuring that disclosures involving harm are escalated appropriately (Dietrich et al., 2023). Furthermore, while robots may facilitate openness, they should serve as complementary tools, not substitutes for human caregivers, ensuring emotional support remains part of a broader care framework (Laban and Cross, 2024). These ethical considerations deserve their own dedicated discussions, which we are currently exploring through parallel ongoing research (Levinson et al., 2024). A key finding from our interviews with subject matter experts as mentioned in (Levinson et al., 2024) is that, while robots should not make judgement calls, they can be a valuable addition to the children’s safeguarding team. However, there is a clear need for interdisciplinary dialogue to thoughtfully inform policy and legal frameworks that define the appropriate roles of robots in children’s lives, particularly in intimate and personal settings (Levinson et al., 2024).
6. Limitations and Future Work
Although we have addressed the research questions defined at the outset of our work, our study has several limitations that we hope to address in the future.
First, we have only considered verbal responses from two tasks of the study. Future work will also focus on verbal responses from other tasks (e.g., Task 1: recall of happy and sad memories, Task 3: picture-based task), children’s drawings, and non-verbal responses (e.g., facial expressions, speech), enabling automatic multimodal robot-led assessment of children’s mental wellbeing. Second, our research is a feasibility study conducted on a non-clinical population to determine whether and how SARs can be used as novel tools to contribute to the holistic understanding of mental health status. In future, we do hope to conduct case-controlled designs in clinical settings with clinical populations through clinician-administered measures. Third, we have not taken into consideration whether/how the varying ethnicities of the children might influence the robotised wellbeing assessment. In future, we aim to investigate how cultural differences might affect the robot-led wellbeing assessment of children.
Fourth, we have not considered whether the children’s past experiences with robots might affect their interaction. In future, we aim to investigate how their attitudes towards robots due to their prior experiences might influence their experience of the robotised wellbeing assessment. Fifth, we have not taken into account wellbeing changes between online form-filling and the experiment sessions. In future, we aim to conduct self-report and parent-report questionnaires concurrently with the robot interaction session to avoid any potential fluctuations in mental wellbeing reporting.
Sixth, the time frame of the longitudinal interactions has been kept relatively focused to 3-6 weeks. While this period allowed us to capture meaningful trends in children’s wellbeing assessments and perceptions of the robot, extending the study over a longer duration could offer additional insights into the sustained dynamics of CRI (Leite et al., 2013). A longer-term study would help to further explore how children’s engagement with the robot evolves over time, especially regarding the potential novelty effect (Smedegaard, 2019). Such a study could determine whether the robot continues to maintain children’s interest and provides sustained benefits, or if engagement declines after additional repeated interactions. Nevertheless, the findings of this study provide an important benchmark, demonstrating that over the 3–6 week period, children’s perceptions of the robot either remained stable or improved, suggesting that the novelty effect did not negatively influence engagement within this timeframe. Future research should seek to replicate and build upon these results by extending interactions to further assess the long-term potential of social robots in child wellbeing assessments.
Additionally, while our study utilized the Nao robot, a widely recognized humanoid platform in HRI (Amirova et al., 2021), it is important to consider that different embodiments of robots may lead to variations in children’s interactions and responses. Prior research has demonstrated that robot embodiment can significantly influence factors such as trust, engagement, and self-disclosure (e.g., Laban et al., 2021; Spitale et al., 2023), potentially affecting wellbeing assessments. While we expect that results would remain consistent with other humanoid robots that support verbal interaction (Vallverdú and Trovato, 2016; Cross and Ramsey, 2021), future research should explore how other embodiments (e.g., more machine-like designs, or animated virtual humans) might evoke different emotional responses or alter children’s perceptions of robots as confidants. Expanding this research to include robots with diverse embodiments will help clarify the generalizability of these findings across various robotic platforms and better inform the role of embodiment in robot-mediated mental health assessments.
@@ 6. Limitations and Future Work
7. Conclusions
Our work has unveiled compelling insights into longitudinally assessing children’s mental wellbeing through online interactions with a robot. We hope that robots can form a valuable addition to the mental wellbeing assessment procedure of children, to aid clinicians and psychologists in identifying cases with potential mental wellbeing concerns so that necessary clinical intervention can be provided in a timely manner. Our study demonstrated that remote robotised wellbeing assessment platforms can effectively gauge children’s wellbeing longitudinally. Furthermore, the user perception of these robot-mediated wellbeing assessment platforms has been sustained or in many instances improved over time. In addition, our exploratory analyses revealed gender-related variability, thus indicating important considerations for developing personalised interactions. As such, incorporating remote robotised platforms in mental wellbeing assessment frameworks offers valuable potential to improve accessibility and support the mental health of the future generation.
@@ Acknowledgments
N. I. Abbasi is supported by the W.D. Armstrong Trust PhD Studentship and the Cambridge Trusts. G. Laban and H. Gunes are supported by the EPSRC project ARoEQ under grant ref. EP/R030782/1. T. Ford and P. Jones acknowledge support from the NIHR ARC East of England and the Cambridge BRC. The equipment used in the study was supported by the University of Cambridge’s Observatory for Human Machine Collaboration (OHMC) Small Equipment Funding. The views expressed are those of the authors and not necessarily those of the NIHR or the Department of Health and Social Care. The authors would also like to thank Miss Amina Itrat Abbasi for her help with the experimental setup.
@@
For the purpose of open access, the authors have applied a Creative Commons Attribution (CC BY 4) licence to any Author Accepted Manuscript version arising.